Find the complete integral of the equation
∂z/∂x[1+(∂z/∂y) ^2]=∂z/∂y(z-a)
1
Expert's answer
2020-10-06T18:28:54-0400

Given differential equation "\frac{∂z}{∂x}[1+(\frac{∂z}{∂y}) ^2]=\frac{∂z}{∂y}(z-a)"


we know that, "p = \frac{∂z}{∂x}" and "q = \frac{∂z}{∂y}"

Let "u = x + by"

Then "p = \frac{dz}{du}" and "q = b\frac{dz}{du}"


Then equation will be

"\frac{dz}{du}[1 +b^2(\frac{dz}{du})^2 ] = b(\frac{dz}{du})(z-a)"


"\frac{dz}{du} = \frac{\sqrt{(bz-ab-1)}}{b}"


Integrating equation,

"\int\frac{bdz}{\sqrt(bz-ab-1)} = \int du"


"2\sqrt{(bz-ab-1)} = u + C"


then

"4(bz-ab-1) = (u+C)^2"

putting value of u,


"4(bz-ab-1) = (x+by+C)^2"

which is the required solution.